Cosmic microwave background radiation fills space with an average energy density of about

4
Answer #1

apply eenrgy density u = 0.5 eo E^2

A. so E^2 = 2* 4*10^-14/8.85*10^-12

E = 0.095 V/m

Erms =


B. I = power/area = Erms^2/uoC

find area from this

as area = pir^2 = power* uoC/Erms^2

pi r^2 = 2000 * 4pi*10^-7*3e8/0.095^2

r = 5156.82 or 5.156 Km =
Answer #2

A-Find the rms value of the electric field associated with this radiation.

=1/2 eE^2= eE^2rms

E rms =sqrt U/e= sqrt 4